OpenStudy (anonymous):

Undefined terms of Geometry: Point= ___dimensions; a ______________________ Line= ___ dimension; a _________ of _____________ Plane=___dimensions; a _______ _____________ Theorems are ______ statements that can be _______________ Postulates are assumed _______ __without _________ I have no Idea what I'm supposed to do please explain so that I may do this as soon as possible.

OpenStudy (anonymous):

@Hero @bohotness @7thGrader @Anon101 @EmogirlAtEmoooocow

OpenStudy (anonymous):

@King.Void. @undeadknight26

OpenStudy (7thgrader):

fcsbhd

OpenStudy (7thgrader):

those are totally the answers lol

OpenStudy (anonymous):

lol

OpenStudy (7thgrader):

ill help u if you help me ok?

OpenStudy (anonymous):

Deal

OpenStudy (7thgrader):

brb

OpenStudy (anonymous):

okayz

OpenStudy (7thgrader):

hich pair of ratios is proportional 7thGrader Medals 0 6/21 and 2/42 25/100 and 50/150 18/24 and 6/12 7/14 and 20/40

OpenStudy (anonymous):

Okay thanx now I'll help you.

OpenStudy (7thgrader):

ok while u find my answer ill find yours

OpenStudy (anonymous):

Lol I gliched out

OpenStudy (7thgrader):

lol

OpenStudy (anonymous):

FUDGE NUGGETS I GLITCHED OUT AGAIN

OpenStudy (7thgrader):

~_~

OpenStudy (anonymous):

right when I was about to post the answer

OpenStudy (7thgrader):

^_~

OpenStudy (7thgrader):

@HailKK I GOT DANCING PIZZA! YAY!

OpenStudy (anonymous):

YAY!!!

OpenStudy (7thgrader):

its gonn taste so good its pepperoni @HailKK

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

\[(6 / 21) + (2 / 42) =0.33333333333\] \[(25 / 100) + (50 / 150) = 0.58333333333\] \[(18 / 24) + (6 / 12) = 1.25\] \[(7 / 14) + (20 / 40) = 1\]

OpenStudy (7thgrader):

so D?

OpenStudy (anonymous):

Those are the answers I cant give you the exact one because of the code of conduct

OpenStudy (7thgrader):

so d?

OpenStudy (7thgrader):

i can't give answers on your question but here hope this helps

OpenStudy (7thgrader):

dudda dudda dudda back down to that site

OpenStudy (anonymous):

Yo @7thGrader get ready for a medal!

OpenStudy (7thgrader):

YAY!

OpenStudy (anonymous):

Just gave it to you

OpenStudy (anonymous):

And I think the answer is D but I would get a second opinion.

OpenStudy (7thgrader):

has t be d